An arrow is fired into the air with an initial velocity of 64 feet per second. The height in
feet of the arrow t seconds after it was shot into the air is given by the function
h(x) = -16t2 + 64t. Find the maximum height of the arrow.

Answer by FrankM(1040)

The form is a parabola opening downward. The maximum height is the vertex.
x=-b/2a = -64/(2*-16) = 2 so the maximum height occurs at t=2 sec.
h=-16(2^2)+64(2) = -64+128 = 64 ft max height.